# miniCTX: Neural Theorem Proving with (Long-)Contexts

### v2

@@ Abstract
Real-world formal theorem proving often depends on a wealth of context, including definitions, lemmas, comments, file structure, and other information. We introduce miniCTX, which tests a model’s ability to prove formal mathematical theorems that depend on new context that is not seen during training. miniCTX contains theorems sourced from real Lean projects and textbooks, each associated with a context that can span tens of thousands of tokens. Models are tasked with proving a theorem given access to code from the theorem’s repository, which contains context that is needed for the proof. As a baseline for miniCTX, we tested fine-tuning and prompting methods that condition theorem proving on preceding context. Both approaches substantially outperform traditional methods that rely solely on state information. We found that this ability to use context is not captured by previous benchmarks such as miniF2F. Alongside miniCTX, we offer ntp-toolkit for automatically extracting and annotating theorem proving data, making it easy to add new projects into miniCTX to ensure that contexts are not seen during training. miniCTX offers a challenging and realistic evaluation of neural theorem provers. [^1] Data and models: HuggingFace. ntp-toolkit: cmu-l3/ntp-toolkit. Evaluation code: cmu-l3/minictx-eval.
@@ 2 Theorem proving with context
We develop miniCTX, a Lean 4 theorem proving benchmark of theorems that depend on newly-defined lemmas, definitions, and proofs from within a project. miniCTX is currently based on 376 theorems from six projects: (1) Prime Number Theorem (PNT) (Kontorovich, 2024a), (2) Polynomial Freiman-Ruzsa Conjecture and its extension with high cross-file dependency (PFR) (Tao, 2023), (3) recent results from the standard mathematical library (Mathlib), (4) an introductory text on theorem proving (HTPI) (Macbeth, 2023), (5) high energy physics formalization in HepLean (HEP) (Tooby-Smith, 2024), and (6) scientific computing formalizations (SciLean) (Skřivan, 2021). Table 2 shows an overview of the dataset. Each theorem in miniCTX consists of the theorem statement, preceding file contents up to the theorem statement, and metadata, in JSON (see §A.1).
@@
      Imported modules, along with separate files recording the premises defined in each imported module,
@@
(tokens) (premises / 100 tokens) (premises / 100 tokens) (lines)
miniF2F Valid/Test 488   153*  3.0
miniCTX PNT 85   10,858 1.87 0.30   3.3
PFR 54   18,059 0.65 1.10 27.2
PFRcross 43   4,351 0.44 2.75   2.7
Mathlib 50   14,440   6.1
HTPI 45   65,082 2.85 0.00 10.7
Hep 61   3,585 5.65 4.25   3.1
SciLean 46   6,249 2.08 9.72   1.8
All 384   18,690 1.94 2.63    8.5
@@
PFR. PFR (Tao, 2023) is a project started in November 2023 that formalizes a proof of the Polynomial Freiman–Ruzsa (PFR) conjecture. We included 54 theorems from PFR. We find that proofs of theorems in PFR tend to be much more monolithic and longer in length than those in Mathlib or other libraries. PFR also defines custom mathematical concepts and notations (such as Ruzsa distance) and a proof typically depends on many lemmas in PFR outside the current file.
PFR_crossfile. PFR_crossfile is an extension of the PFR split, which includes additional problems to further evaluate cross-file dependencies. To evaluate models’ performance on problems with extensive cross-file dependencies, we added 43 theorems from the files Entropy.Group and Entropy.Kernel.Group, which have the highest cross-file dependencies in the project. These problems contain three times the number of cross-file premises compared to other math splits, making them a strong candidate for evaluating a model’s ability to utilize cross-file premises effectively.
Recent Mathlib Commits. Mathlib (Mathlib Community, 2020), is Lean’s largest community-maintained repository, encompassing a wide range of mathematical concepts, programming APIs, and common tactics. It is commonly used for training theorem-proving models and is frequently updated with new definitions, theorems, and refactorings. Therefore, to avoid data contamination, we included 50 theorems newly added to Mathlib in April 2024, by filtering recent Mathlib commits to ones that only add new theorems. Assuming that the model was trained prior to April 2024, the Mathlib split guarantees the evaluation of theorem-level generalization.
@@
Recent SciLean Commits. SciLean (Skřivan, 2021) is an open-source project designed to formalize concepts in scientific computing using Lean 4. SciLean aims to provide formalized tools and frameworks for efficiently representing and proving properties of numerical methods, differential equations, and optimization problems, bridging applied mathematics and computer science. Given Scilean has been created for three years and accessed by multiple projects, similar to Mathlib split, we selected 46 theorems added to Scilean in April 2024 in order to ensure the testing data is more likely unseen for evaluating models. The newly added problems are mostly supplements to existing files, providing an opportunity to evaluate the model’s ability to apply and learn from previous lemmas.
@@ 3.2 Problem selection methodology
    Recency-based selection: For popular libraries such as Mathlib and SciLean, which are highly likely to be used for training, we aim to select newly added theorems. This is achieved by sorting theorems based on the timestamp of when the theorem was first added to the project, which is extracted by ntp-toolkit. This strategy helps mitigate data contamination risks.
@@
    Dependency-based selection: To explicitly evaluate models’ performance in context-dependent proving, we selected files based on the in-file and cross-file premise labels available in our benchmark. For the Prime split, we chose the file with the highest number of in-file dependencies, while for PFRcrossfile, we selected files with the highest cross-file dependencies. For HEPLean, we selected files with a balance of both types. The level of dependency is also extracted automatically by ntp-toolkit.
@@ 3.3 Key features and challenges
Real-world theorem proving. Unlike popular benchmarks (e.g., miniF2F (Zheng et al., 2022), ProofNet (Azerbayev et al., 2023), FIMO (Liu et al., 2023)) that focus on isolated competition problems, real-world research-level theorem proving is heavily dependent on rich mathematical contexts. Therefore, miniCTX includes real-world, complex theorems from a variety of ongoing Lean projects, such as Prime Number Theorem (Prime) and Polynomial Freiman-Ruzsa Conjecture (PFR). They rigorously test a model’s ability in real-world formalization projects. This diversity contrasts with the LeanDojo benchmark (Yang et al., 2023), which focuses solely on Mathlib, enabling miniCTX to better test a model’s generalization in different settings.
@@
Automatically updating the benchmark. Most modern neural theorem provers use a large language model as a backbone. Therefore, it is crucial to ensure that evaluation content is not seen during (pre-)training, a problem not addressed by previous benchmarks. miniCTX’s format is amenable to periodically updating the benchmark with new projects to ensure that proofs are not seen by language models trained prior to a particular date. Future periodic updates will be automatically extracted from new Lean projects and commits using ntp-toolkit (§3.4). See Figure 2 for an illustration.
@@ 4.1 Baselines
File-tuning. A drawback to the setups above is that they only consider static (state, tactic) pairs, and do not take into account new context, including comments, definitions, lemmas, and file structure, encountered during test time. Therefore, we test whether supplying context $c$, in the form of preceding file contents, to the model improves performance. Similar to state-tactic tuning, we fine-tune a 1.3b model on human-written (state $x_{t}$, context $c$, tactic $y_{t}$) triples to generate tactics based on the current context and proof state, resulting in the file-tuned model (see Figure 3 for an illustration).
@@
Table 3: Performance comparison (%) of different models on miniF2F and miniCTX.
miniF2F miniCTX
@@ 4.2 Results
Evaluation on miniF2F. We evaluate baselines on miniF2F, a standard benchmark based on competition problems that do not require context. The file-tuned model improves very little beyond the state-tactic model (33.61% vs. 32.79%), showing that the dramatic difference in context-dependent proving abilities seen on miniCTX cannot be captured by miniF2F.
@@ D.1 Dependency distribution
Figure 5: Percentage of different dependencies in the human-written proof of theorems in miniCTX
@@ Appendix E Dataset hosting and maintenance
miniCTX is released on HuggingFace: l3lab/miniCTX, distributed under the Apache 2.0 license. Data extraction tool ntp-toolkit is released on GitHub: cmu-l3/ntp-toolkit, under the MIT license. We note that the underlying data for the individual splits of miniCTX are also released under the Apache 2.0 license. We include the licensing information in the dataset repository. We plan to regularly update and maintain the dataset to include examples from new projects.
@@ F.1 Preliminary
The evaluation code relies heavily on the Lean REPL (Lean Prover Community, 2024), which operates within the project environment. Therefore, it is essential that the project builds without any errors. Additionally, the version of Lean used in the project should match the version supported by the REPL. While the Lean REPL supports versions $\geq$ 4.3.0, for the best experience with data extraction and evaluation, we recommend evaluating projects that use Lean version 4.7.0 (all miniCTX theorems are in 4.7.0). We plan to continuously update ntp-toolkit to support newer versions.
@@ F.2 Using the ntp-toolkit
    Installation: Clone the ntp-toolkit repository from GitHub to your local machine, and checkout the Lean version tag corresponding to the extracted project (e.g., v4.7.0). Ensure that you have the required dependencies installed, as listed in the repository’s README file.

### v3

@@ Abstract
Real-world formal theorem proving often depends on a wealth of context, including definitions, lemmas, comments, file structure, and other information. We introduce miniCTX, which tests a model’s ability to prove formal mathematical theorems that depend on new context that is not seen during training. miniCTX contains theorems sourced from real Lean projects and textbooks, each associated with a context that can span tens of thousands of tokens. Models are tasked with proving a theorem given access to code from the theorem’s repository, which contains context that is needed for the proof. As a baseline for miniCTX, we tested fine-tuning and prompting methods that condition theorem proving on preceding context. Both approaches substantially outperform traditional methods that rely solely on state information. We found that this ability to use context is not captured by previous benchmarks such as miniF2F. Alongside miniCTX, we offer ntp-toolkit for automatically extracting and annotating theorem proving data, making it easy to add new projects into miniCTX to ensure that contexts are not seen during training. miniCTX offers a challenging and realistic evaluation of neural theorem provers.111 Project page: https://cmu-l3.github.io/minictx. Please refer to our project page for our dataset and evaluation links, and future updates including miniCTX-v2.
@@ 2 Theorem proving with context
We develop miniCTX, a Lean 4 theorem proving benchmark of theorems that depend on newly-defined lemmas, definitions, and proofs from within a project. miniCTX is currently based on 762 theorems from six projects: (1) Prime Number Theorem (PNT) (Kontorovich, 2024a), (2) Polynomial Freiman-Ruzsa Conjecture and its extension with high cross-file dependency (PFR) (Tao, 2023), (3) recent results from the standard mathematical library (Mathlib), (4) an introductory text on theorem proving (HTPI) (Macbeth, 2023), (5) high energy physics formalization in HepLean (HEP) (Tooby-Smith, 2024), and (6) scientific computing formalizations (SciLean) (Skřivan, 2021). These theorems are equally split into 381 validation and 381 test theorems. Table 2 shows an overview of the dataset. Each theorem in miniCTX consists of the theorem statement, preceding file contents up to the theorem statement, and metadata, in JSON (see §A.1).
@@
      Imported modules (for premise selection support),
@@
valid + test (tokens) (premises / 100 tokens) (premises / 100 tokens) (lines)
miniF2F 244 + 244   153*  3.0
miniCTX PNT 85 + 85   10,858 1.87 0.30   3.3
PFR 51 + 51   18,059 0.65 1.10 27.2
PFRcross 43 + 43   04,351 0.44 2.75   2.7
Mathlib 50 + 50   14,440   6.1
HTPI 45 + 45   65,082 2.85 0.00  10.7
HEP 61 + 61   03,585 5.65 4.25   3.1
SciLean 46 + 46   06,249 2.08 9.72   1.8
All 381 + 381   18,690 1.94 2.63    8.5
@@
PFR. PFR (Tao, 2023) is a project started in November 2023 that formalizes a proof of the Polynomial Freiman–Ruzsa (PFR) conjecture. We included 51 validation and 51 test theorems from PFR. We find that proofs of theorems in PFR tend to be much more monolithic and longer in length than those in Mathlib or other libraries. PFR also defines custom mathematical concepts and notations (such as Ruzsa distance) and a proof typically depends on many lemmas in PFR outside the current file.
PFRcrossfile. PFRcrossfile is an extension of the PFR split, which includes additional problems to further evaluate cross-file dependencies. To evaluate models’ performance on problems with extensive cross-file dependencies, we added 43 test theorems from Entropy.Group and Entropy.Kernel.Group, which have the highest cross-file dependencies in the project. These problems contain three times the number of cross-file premises compared to other math splits, making them a strong candidate for evaluating a model’s ability to utilize cross-file premises. 43 validation theorems are chosen similarly.
Recent Mathlib Commits. Mathlib (Mathlib Community, 2020), is Lean’s largest community-maintained repository, encompassing a wide range of mathematical concepts, programming APIs, and common tactics. It is commonly used for training theorem-proving models and is frequently updated with new definitions, theorems, and refactorings. Therefore, to avoid data contamination, we included 50 test and 50 validation theorems newly added to Mathlib since March 2024, by filtering recent Mathlib commits to ones that only add new theorems. Assuming that the model was trained prior to April 2024, the Mathlib split guarantees the evaluation of theorem-level generalization.
@@
Recent SciLean Commits. SciLean (Skřivan, 2021) is an open-source project designed to formalize concepts in scientific computing using Lean 4. SciLean aims to provide formalized tools and frameworks for efficiently representing and proving properties of numerical methods, differential equations, and optimization problems, bridging applied mathematics and computer science. Given Scilean has been created for 3 years and accessed by multiple projects, similar to our Mathlib split, we selected 46 test and 46 validation theorems added to SciLean since March 2024 in order to ensure the data is more likely unseen for evaluating models. The newly added problems are mostly supplements to existing files, so they evaluate the model’s ability to apply and learn from previous lemmas.
@@ 3.2 Problem selection methodology
    Recency-based selection: For popular libraries such as Mathlib and SciLean, which are highly likely to be used for training, we aim to select newly added theorems. This is achieved by sorting theorems based on the timestamp of when the theorem was first added to the project, which is extracted by ntp-toolkit. This helps mitigate data contamination.
@@
    Dependency-based selection: To explicitly evaluate models’ performance in context-dependent proving, we selected files based on the in-file and cross-file premise labels available in our benchmark. For the PNT split, we chose the file with the highest number of in-file dependencies, while for PFRcrossfile, we selected files with the highest cross-file dependencies. For HEP, we selected files with a balance of both types. The level of dependency is also extracted automatically by ntp-toolkit.
@@ 3.3 Key features and challenges
Real-world theorem proving. Unlike popular benchmarks (e.g., miniF2F (Zheng et al., 2022), ProofNet (Azerbayev et al., 2023), FIMO (Liu et al., 2023)) that focus on isolated competition problems, real-world research-level theorem proving is heavily dependent on rich mathematical contexts. Therefore, miniCTX includes real-world, complex theorems from a variety of ongoing Lean projects, such as Prime Number Theorem (PNT) and Polynomial Freiman–Ruzsa Conjecture (PFR). They rigorously test a model’s ability in real-world formalization projects. This diversity contrasts with the LeanDojo benchmark (Yang et al., 2023), which focuses solely on Mathlib, enabling miniCTX to better test a model’s generalization in different settings.
@@
Automatically updating the benchmark. Most modern neural theorem provers use a large language model as a backbone. Therefore, it is crucial to ensure that evaluation content is not seen during (pre-)training, a problem not addressed by previous benchmarks (see §G). We plan to update miniCTX periodically to include theorems beyond a certain cut-off date (see our project page). Future updates will be automatically extracted from new Lean projects using ntp-toolkit (§3.4). See Figure 2.
@@ 4.1 Baselines
File-tuning. A drawback to the setups above is that they only consider static (state, tactic) pairs, and do not take into account new context, including comments, definitions, lemmas, and file structure, encountered during test time. Therefore, we test whether supplying context $c$, in the form of preceding file contents, to the model improves performance. Similar to state-tactic tuning, we fine-tune DeepSeek-Coder-1.3b on human-written (state $x_{t}$, context $c$, tactic $y_{t}$) triples to generate tactics based on the current context and proof state, resulting in the file-tuned model (see Figure 3).
@@
Table 3: Performance comparison (%) of different models on miniF2F-test and miniCTX-test.
miniF2F miniCTX-test
@@ 4.2 Results
Evaluation on miniF2F. We evaluate baselines on miniF2F, a standard benchmark based on competition problems that do not require context. We use import statements as context for the file-tuned model. The file-tuned model improves very little beyond the state-tactic model (33.61% vs. 32.79%), showing that the dramatic difference in context-dependent proving abilities seen on miniCTX cannot be captured by miniF2F.
@@ Appendix D Additional results and analysis
We present additional analysis on the composition of miniCTX and additional quantitative and qualitative evaluation of performance on miniCTX. All tests in this section are done on miniCTX-test.
@@ D.1 Dependency distribution
Figure 5: Percentage of different dependencies in the human-written proof of theorems in miniCTX.
@@ Appendix E Dataset hosting and maintenance
miniCTX is released on HuggingFace: l3lab/miniCTX, distributed under the Apache 2.0 license. Data extraction tool ntp-toolkit is released on GitHub: cmu-l3/ntp-toolkit, under the MIT license. We note that the underlying data for the individual splits of miniCTX are also released under the Apache 2.0 license. We include the licensing information in the dataset repository. We plan to regularly update and maintain the dataset to include examples from new projects. For information about future updates such as miniCTX-v2, please refer to our project page: https://cmu-l3.github.io/minictx.
@@ F.1 Preliminary
The evaluation code relies heavily on the Lean REPL (Lean Prover Community, 2024), which operates within the project environment. Therefore, it is essential that the project builds without any errors. Additionally, the version of Lean used in the project should match the version supported by the REPL. While the Lean REPL supports versions $\geq$ 4.3.0, for the best experience with data extraction and evaluation, we recommend evaluating projects that use Lean version 4.7.0 or higher (all miniCTX theorems are in 4.7.0). We plan to continuously update ntp-toolkit to support newer versions.
@@ F.2 Using the ntp-toolkit
    Installation: Clone the ntp-toolkit repository from GitHub to your local machine. Currently, to use ntp-toolkit for Lean version 4.7, switch to the lean-v4.7.0 branch; for 4.8.0 or above, use the main branch. Ensure that you have the required dependencies installed, as listed in the repository’s README file.
@@
Appendix G Data contamination in existing benchmarks
One of the contributions of miniCTX is the temporal split: using ntp-toolkit, miniCTX only includes theorems created after a certain cut-off date. This ensures LLMs trained before this date have not seen problems in miniCTX. We claimed that existing benchmarks face significant risks of data contamination, and here we provide some evidence to this claim.
With regards to contamination, existing benchmarks can be largely categorized as either extracted from a standard mathematical library (e.g. Mathlib), or curated as competition-style problems. For the former category, evaluation benchmarks include Mathlib test problems extracted by LeanDojo (Yang et al., 2023) or PACT (Han et al., 2022), as well as datasets like CoqGym (Yang & Deng, 2019) and PISA (Jiang et al., 2021) in other formal languages. They are inherently at risk of contamination, because they are statically sourced from projects on GitHub, many of which have existed for years (e.g. Mathlib has existed since 2017). On the other hand, virtually all modern language models are pre-trained on public GitHub code, and have therefore seen the proofs. This partially invalidates any evaluation results using these benchmarks.
On the other hand, we have empirically found many solutions of datasets like miniF2F (Zheng et al., 2022) online. The original repository of miniF2F already contains full solutions to 72 of the 244 test problems (in Lean 3, which is directly translatable to Lean 4) [^3] https://github.com/openai/miniF2F. Moreover, of the remaining problems, 12 IMO problems have full solutions in Mathlib [^4] https://github.com/leanprover-community/mathlib4/tree/master/Archive/Imo and Compfiles (a catalog of Lean proofs of competition problems) [^5] https://github.com/dwrensha/compfiles, at the time of writing. In total, at least 84 of the 244 problems in miniF2F-test, including 12 of the 19 IMO problems, have full solutions on GitHub. This may be an inherent issue of small competition-style benchmarks, as the theorem-proving community itself has significant interest in formalizing competition problems and solutions. The scarcity of problems from competitions like IMO each year, as well as the bias to only formalizing some categories (geometry problems are often not formalized due to difficulty) also contribute to this issue. Whether temporal split and other mitigation methods can be applied to competition-style benchmarks may also be an interesting future direction of research.